Hello out there,I need some help with some problems,the first one is,Simplify:6(S.ROOT)7-(S.ROOT)7.The next one is,Simplify:2(S.ROOT)5+2(S.ROOT)5.The third one is,Simplify:(S.ROOT)100Y^2 Z^4.And the last one is,Simplify:(S.ROOT)50x^2,thanks to whoever can help.
Answer by edjones(7311)

A) 6sqrt(7)-sqrt(7)
substitute a for sqrt(7)
6a-a=???
sure! 5a
now just substitute sqrt(7) for a.
5sqrt(7)
.
B) you can do this.
.
C) sqrt(100y^2z^4)=10yz^2
.
D) sqrt(50x^2)=x sqrt(25*2)=5x sqrt(2)
.
Ed